# Combinatorial invariance for the coefficient of $q$ in Kazhdan–Lusztig polynomials

### v1

@@ Abstract.
We prove the combinatorial invariance of the coefficient of $q$ in Kazhdan–Lusztig polynomials for arbitrary Coxeter groups. As a result, we obtain the Combinatorial Invariance Conjecture for Bruhat intervals of length at most $6$. We also prove the Gabber–Joseph conjecture for the second-highest $\operatorname{Ext}$ group of a pair of Verma modules, as well as the combinatorial invariance of the dimension of this group.
@@ 1.1. The Combinatorial Invariance Conjecture
The Kazhdan–Lusztig polynomials $P_{u,v}(q)\in\mathbb{N}[q]$, indexed by elements $u,v$ in a Coxeter group $W$, underlie deep connections between the singularities of Schubert varieties, representations of Lie algebras, bases for Hecke algebras, and Soergel bimodules [KL79, EW14]. Yet the Combinatorial Invariance Conjecture, due to Lusztig (c. 1983) and to Dyer [Dye87], asserts that they depend only on the Bruhat interval $[u,v]$ as an abstract poset.
@@
Let $u\leq v\in W$ and $u^{\prime}\leq v^{\prime}\in W^{\prime}$ be such that $u,v]\cong[u^{\prime},v^{\prime}]$. Then $q]P_{u,v}=[q]P_{u^{\prime},v^{\prime}}$, $q]R_{u,v}=[q]R_{u^{\prime},v^{\prime}}$, $q^{\ell(u,v)-1}]R_{u,v}=[q^{\ell(u,v)-1}]R_{u^{\prime},v^{\prime}}$, and $q^{\ell(u,v)}-2]\widetilde{R}_{u,v}=[q^{\ell(u^{\prime},v^{\prime})}-2]\widetilde{R}_{u^{\prime},v^{\prime}}$.
@@
Table 1. Coefficients of the $P$-, $R$-, and $\widetilde{R}-$polynomials which are known to be combinatorially invariant for all Coxeter groups. Prior to the current work, the most recent update to this table was in 1987.
@@
Recent results [CGG+25, GLSBss, BY25] have established that the coordinate ring of a Richardson variety is a cluster algebra [FZ02]. Although we do not need this fact, it is interesting to note that, when $W$ is crystallographic, the specific diamond-generating set that we construct in the proof of Theorem 1.2 corresponds to the frozen variables in this cluster structure. Thus we also obtain the combinatorial invariance of the numbers of frozen and of mutable variables in this cluster structure.
@@ 1.2. Extensions of Verma modules
Let $\mathfrak{g}$ be a complex semisimple Lie algebra with Borel subalgebra $\mathfrak{b}$ and Weyl group $W$. For $w\in W$, let $M_{w}$ be the Verma module of highest weight $w\lambda-\rho$, where $\lambda$ is any anti-dominant regular integral weight and $\rho$ is the half-sum of the positive roots. Inspired by [GJ81], the Gabber–Joseph conjecture [^1] Despite this standard name, the conjecture does not actually appear in the work of Gabber–Joseph. asserted that for $u\leq v$ the polynomial
@@ 1.3. Outline
In Section 2 we recall background on Kazhdan–Lusztig polynomials, Bruhat graphs and reflection orders, and Richardson varieties. In Section 3 we introduce divergences of increasing paths and prove Theorem 1.2. In Section 4 we combine this with some geometric arguments to prove Theorem 1.6.
@@ 2.2. Reflection orders and the Bruhat graph
An initial section of a reflection order is a set $X\subseteq\Phi^{+}$ so that if $\beta\in X$ and $\alpha\in\Phi^{+}$ with $\alpha\prec\beta$, then $\alpha\in X$. An inversion set is a set of roots of the form $N(w)\coloneqq\{\beta\in\Phi^{+}\mid w^{-1}\beta\not\in\Phi^{+}\}$ for some $w\in W$. If $s\in S$ and $\ell(w)<\ell(ws)$, then $N(ws)=N(w)\sqcup\{w\alpha_{s}\}$. In this way, a reduced expression $w=s_{i_{1}}\cdots s_{i_{k}}$ is equivalent to a chain of inversion sets $\varnothing\subset N(s_{i_{1}})\subset N(s_{i_{1}}s_{i_{2}})\subset\cdots\subset N(w)$.
@@ 3. Divergences of increasing paths
Figure 3. The inductive construction of a supporting chain. The action of $s$ is indicated with blue edges. The second-length path for $[su,sv]$ and supporting chain $\mathcal{C}^{\prime}=\{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}sx_{1},\circ}\}$ are shown with dashed orange and purple lines, respectively. The supporting chain for $[u,v]$ is $\mathcal{C}=\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{1},\bullet,sy}\}$, shown in red.
@@
Write the elements of $\mathcal{C}^{\prime}$ in order as $c_{0},\ldots,c_{k}$. Assume to the contrary that there exist $c_{i},c_{i+1}$ such that $sc_{i}>c_{i}$ and $sc_{i+1}<c_{i+1}$. This would imply, by the lifting property of Bruhat order, that $sc_{i}=c_{i}$. Moreover, since $c_{i}$ covers $sy_{i}$, the lifting property further implies that $s\cdot sy_{i}<sy_{i}$, which is a contradiction. Hence there can be no $i$ such that $sc_{i}<c_{i}$. From this we can see that $\mathcal{C}$ is a saturated chain, since $\mathcal{C}^{\prime}\sqcup\{y_{j+1},\ldots,y_{r}\}$ is a saturated chain, each of whose elements is taken upward by $s$. Moreover $sy_{r}\leq z$, by using the lifting property with the fact that $sz<z$. Hence $\mathcal{C}$ is a supporting chain. ∎
@@ 4. Application to Gabber–Joseph coefficients
Proposition 4.1 can also be deduced from properties of the mixed Hodge structures of cluster varieties [LS22, GLSBss, CGG+25].

### v2

@@ Abstract.
We prove the combinatorial invariance of the coefficient of $q$ in Kazhdan–Lusztig polynomials for arbitrary Coxeter groups. As a result, we obtain the Combinatorial Invariance Conjecture, of Lusztig and of Dyer, also for Bruhat intervals of length at most $6$. We also prove the Gabber–Joseph conjecture for the second-highest $\operatorname{Ext}$ group of a pair of Verma modules, as well as the combinatorial invariance of the dimension of this group, and of the numbers of frozen and of mutable variables in the cluster structure on Richardson varieties.
@@ 1.1. The Combinatorial Invariance Conjecture
The Kazhdan–Lusztig polynomials $P_{u,v}(q)\in\mathbb{N}[q]$, indexed by elements $u,v$ in a Coxeter group $W$, underlie deep connections between the singularities of Schubert varieties, representations of Lie algebras, bases for Hecke algebras, and Soergel bimodules [KL79, EW14]. Yet the Combinatorial Invariance Conjecture, due to Lusztig (c. 1983) and to Dyer [Dye87], asserts that they depend only on the Bruhat interval $[u,v]$ as an abstract poset (a property we will more generally refer to as combinatorial invariance).
@@
Let $u\leq v\in W$ and $u^{\prime}\leq v^{\prime}\in W^{\prime}$ be such that $u,v]\cong[u^{\prime},v^{\prime}]$. Then $q]P_{u,v}=[q]P_{u^{\prime},v^{\prime}}$, $q]R_{u,v}=[q]R_{u^{\prime},v^{\prime}}$, $q^{\ell(u,v)-1}]R_{u,v}=[q^{\ell(u^{\prime},v^{\prime})-1}]R_{u^{\prime},v^{\prime}}$, and $q^{\ell(u,v)}-2]\widetilde{R}_{u,v}=[q^{\ell(u^{\prime},v^{\prime})}-2]\widetilde{R}_{u^{\prime},v^{\prime}}$.
@@
Table 1. Coefficients of the $P$-, $R$-, and $\widetilde{R}-$polynomials which are known to be combinatorially invariant for all Coxeter groups.
@@
Recent results [CGG+25, GLSBS22, GLSB25] have established that the coordinate ring of a Richardson variety is a cluster algebra [FZ02]. Although we do not need this fact, it is interesting to note that, when $W$ is a finite Weyl group, the specific diamond-generating set that we construct in the proof of Theorem 1.2 corresponds to the frozen variables in this cluster structure. Thus we also obtain the combinatorial invariance of the numbers of frozen and of mutable variables in this cluster structure (see Section 5).
@@ 1.2. Extensions of Verma modules
Let $\mathfrak{g}$ be a complex semisimple Lie algebra with Borel subalgebra $\mathfrak{b}$ and Weyl group $W$. For $w\in W$, let $M_{w}$ be the Verma module of highest weight $w\lambda-\rho$, where $\lambda$ is any anti-dominant regular integral weight and $\rho$ is the half-sum of the positive roots. Inspired by [GJ81], the Gabber–Joseph conjecture [^1] Despite this standard name, the conjecture does not actually appear in the work of Gabber–Joseph which inspired it. asserted that for $u\leq v$ the polynomial
@@
We prove Theorem 1.6 by establishing new vanishing results for the mixed Hodge structure of open Richardson varieties (see Proposition 4.1), which may be of independent interest.
@@ 1.3. Outline
In Section 2 we recall background on Kazhdan–Lusztig polynomials, Bruhat graphs and reflection orders, and Richardson varieties. In Section 3 we introduce divergences of increasing paths and prove Theorem 1.2. In Section 4 we combine this with some geometric arguments to prove Theorem 1.6. In Section 5 we relate our results to the cluster structure on Richardson varieties, and derive some corollaries.
@@ 2.2. Reflection orders and the Bruhat graph
An initial section of a reflection order is a set $X\subseteq\Phi^{+}$ so that if $\beta\in X$ and $\alpha\in\Phi^{+}$ with $\alpha\prec\beta$, then $\alpha\in X$. An inversion set is a set of roots of the form $N(w)\coloneqq\{\beta\in\Phi^{+}\mid w^{-1}\beta\not\in\Phi^{+}\}$ for some $w\in W$. If $s\in S$ and $\ell(w)<\ell(ws)$, then $N(ws)=N(w)\sqcup\{w\alpha_{s}\}$. In this way, a reduced expression $w=s_{1}\cdots s_{\ell}$ is equivalent to a chain of inversion sets $\varnothing\subset N(s_{1})\subset N(s_{1}s_{2})\subset\cdots\subset N(w)$.
@@ 3. Divergences of increasing paths
Figure 3. The inductive construction of a supporting chain. The action of $s$ is indicated with blue edges. The second-length path for $[su,sv]$ is shown with dashed orange lines and the supporting chain $\mathcal{C}^{\prime}=\{{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}sx_{1},\circ}\}$ is shown in purple. The supporting chain for $[u,v]$ is $\mathcal{C}=\{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}x_{1},\bullet,sy}\}$, shown in red.
@@
Write the elements of $\mathcal{C}^{\prime}$ in order as $c_{0},\ldots,c_{k}$. Assume to the contrary that there exist $c_{i},c_{i+1}$ such that $sc_{i}>c_{i}$ and $sc_{i+1}<c_{i+1}$. This would imply, by the lifting property of Bruhat order, that $sc_{i}=c_{i+1}$. Moreover, since $c_{i}$ covers $sy_{i}$, the lifting property further implies that $s\cdot sy_{i}<sy_{i}$, which is a contradiction. Hence there can be no $i$ such that $sc_{i}<c_{i}$. From this we can see that $\mathcal{C}$ is a saturated chain, since $\mathcal{C}^{\prime}\sqcup\{y_{j+1},\ldots,y_{r}\}$ is a saturated chain, each of whose elements is taken upward by $s$. Moreover $sy_{r}\leq z$, by using the lifting property with the fact that $sz<z$. Hence $\mathcal{C}$ is a supporting chain. ∎
@@ 4. Application to Gabber–Joseph coefficients
Proposition 4.1 can also be deduced using properties of the mixed Hodge structures of cluster varieties [LS22] (Richardson varieties are known to be cluster varieties by [GLSB25, CGG+25]).
@@
5. Connection to cluster structure on Richardson varieties
In this section, we briefly describe the features of the cluster structure on a Richardson variety $\mathcal{R}_{u,v}$ which motivated the definition of the diamond-generating set $F_{u,v}$ from Theorem 3.5 and derive some consequences for the cluster structure. The proof of the theorem, however, does not depend on this connection. Indeed Theorem 3.5 applies in arbitrary Coxeter groups, while Richardson varieties exist only in crystallographic type, and are known to be cluster varieties only for finite Weyl groups (see [BY25] for an upper cluster structure in crystallographic type).
Let $W$ be the Weyl group of a complex semisimple Lie group $G$, and let $u,v\in W$. We write $\mathbb{C}[\mathcal{R}_{u,v}]$ for the coordinate ring of the Richardson variety $\mathcal{R}_{u,v}$. A cluster structure on $\mathcal{R}_{u,v}$ includes distinguished elements of $\mathbb{C}[\mathcal{R}_{u,v}]$, called cluster variables, together with distinguished sets of cluster variables called clusters [^2] In addition to cluster variables and clusters, there is more data required to give a cluster structure, but we shall not discuss it here. See [FZ02] for a definition of cluster algebra. . Each cluster contains $\dim\mathcal{R}_{u,v}=\ell(u,v)$ cluster variables. The cluster variables which are in every cluster are called frozen variables. The remaining cluster variables are called mutable variables.
Proposition 5.1.
Let $u\leq v$ be elements of a finite Weyl group $W$. The diamond-generating set $F_{u,v}$ is in bijection with the set of frozen variables for the cluster structure on $\mathbb{C}[\mathcal{R}_{u,v}]$. Consequently, there are $d_{u,v}$ frozen and $\ell(u,v)-d_{u,v}$ mutable variables in each cluster, and the number of these is a combinatorial invariant.
Proof.
The bijection is constructed in the remainder of this section. As noted in the proof of Theorem 3.5, we have $|F_{u,v}|=d_{u,v}$. This quantity is combinatorially invariant by Theorem 1.2. ∎
In order to construct the desired bijection, we now discuss the specifics of the cluster structure introduced in [GLSB25]. Let $\mathbf{v}=s_{1}\cdots s_{\ell(v)}$ be a reduced expression for $v$. Then loc. cit. defines a cluster associated to $\mathbf{v}$, which we denote $X_{\mathbf{v}}$. To index the variables in this cluster, they use the notion of a subexpression for $u$. We write $\mathbf{v}_{(k)}=s_{1}\cdots s_{k}$ for the $k$th partial product of $\mathbf{v}$. We will similarly describe subexpressions of $\mathbf{v}$ by their partial products; hence a subexpression corresponds to a sequence $\mathbf{u}_{(k)}$ such that $\mathbf{u}_{(0)}=e$, $\mathbf{u}_{(\ell(v))}=u$, and, for all $1\leq k\leq\ell(v)$, $\mathbf{u}_{(k)}=\mathbf{u}_{(k-1)}s_{k}$ or $\mathbf{u}_{(k)}=\mathbf{u}_{(k-1)}$.
Definition 5.2.
The positive subexpression of $\mathbf{v}$ is the subexpression $\mathbf{u}^{+}$ whose partial products are defined recursively by
$\mathbf{u}^{+}_{(\ell(v))}=u$
$\mathbf{u}^{+}_{(k-1)}=\min(\mathbf{u}^{+}_{(k)},\mathbf{u}^{+}_{(k)}s_{k}).$
Given a subexpression $\mathbf{u}$, we set $J_{\mathbf{u}}\coloneqq\{k\in[\ell(v)]\mid\mathbf{u}_{(k)}=\mathbf{u}_{(k-1)}\}.$
The cluster variables in $X_{\mathbf{v}}$ are indexed by $J_{\mathbf{u}^{+}}$. Among these, the mutable variables correspond to certain “almost positive subexpressions”.
Definition 5.3.
For $a\in J_{\mathbf{u}^{+}}$, let $\mathbf{u}^{\langle a\rangle}_{(\ell(v))}\coloneqq u$ and define recursively
$\mathbf{u}^{\langle a\rangle}_{(k-1)}\coloneqq\begin{cases}\max(\mathbf{u}_{(k)}^{\langle a\rangle},\mathbf{u}^{\langle a\rangle}_{(k)}s_{k})&\text{if $k=a$}\\ \min(\mathbf{u}_{(k)}^{\langle a\rangle},\mathbf{u}^{\langle a\rangle}_{(k)}s_{k})&\text{if $k\neq a$}.\end{cases}$
We say $a\in J_{\mathbf{u}^{+}}$ is mutable if $\mathbf{u}^{\langle a\rangle}_{(0)}=e$, otherwise $a$ is frozen. If $a$ is mutable, then $\mathbf{u}^{\langle a\rangle}$ is called an almost positive subexpression of $\mathbf{v}$.
We now describe how to relate subexpressions of $\mathbf{v}$ to paths in $\widehat{\Gamma}$.
Definition 5.4.
Let $\mathbf{u}$ be a subexpression of $\mathbf{v}$. Let the indices in $J_{\mathbf{u}}$ be $i_{1}<\cdots<i_{r}$. Define $\operatorname{path}(\mathbf{u})$ to be the path $x_{0}-x_{1}-\cdots-x_{r}$ in $\widehat{\Gamma}$ with vertex sequence
$u^{-1}-v^{-1}\mathbf{v}_{(i_{r}-1)}\mathbf{u}_{(i_{r}-1)}^{-1}-v^{-1}\mathbf{v}_{(i_{r-1}-1)}\mathbf{u}_{(i_{r-1}-1)}^{-1}-\cdots-v^{-1}\mathbf{v}_{(i_{1}-1)}\mathbf{u}_{(i_{1}-1)}^{-1}=v^{-1}.$
We will be particularly interested in the case where $\operatorname{path}(\mathbf{u})$ is in fact a path in the directed Bruhat graph $\Gamma$. The following notion is due to Deodhar [Deo85].
Definition 5.5.
A subexpression $\mathbf{u}$ of $\mathbf{v}$ is distinguished if, for all $k$ such that $\mathbf{u}_{(k-1)}s_{k}<\mathbf{u}_{(k-1)}$, we have $\mathbf{u}_{(k)}=\mathbf{u}_{(k-1)}s_{k}$.
The positive and almost positive subexpressions of $\mathbf{v}$ are always distinguished. The positive subexpression is the unique distinguished subexpression $\mathbf{u}$ with $|J_{\mathbf{u}}|=\ell(u,v)$, and the almost positive subexpressions are exactly the subexpressions $\mathbf{u}$ with $|J_{\mathbf{u}}|=\ell(u,v)-2$.
To describe which paths arise as $\operatorname{path}(\mathbf{u})$ for a distinguished subexpression $\mathbf{u}$, we use reflection orders. A reflection order $\prec$ is said to be compatible with the reduced expression $\mathbf{v}$ if $N(v^{-1}\mathbf{v}_{(i)})$ is an initial section of $\prec$ for all $0\leq i\leq\ell(v)$. Equivalently, the smallest $\ell(v)$ positive roots are
$\Phi_{1}(\mathbf{v})\prec\Phi_{2}(\mathbf{v})\prec\cdots\prec\Phi_{\ell(v)-1}(\mathbf{v})\prec\Phi_{\ell(v)}(\mathbf{v}),$
where $\Phi_{j}(\mathbf{v})$ denotes the root $v^{-1}\mathbf{v}_{(\ell(v)-j+1)}\alpha_{s_{\ell(v)-j+1}}$, the unique root $\beta$ with
(2) $\mathbf{v}_{(\ell(v)-j)}^{-1}vt_{\beta}=\mathbf{v}^{-1}_{(\ell(v)-j+1)}v.$
13241342142323143124143223412413314232142431324134123421
Figure 4. The Bruhat graph for $[1324,3421]$ in $S_{4}$. The longest increasing path $\gamma_{0}$ is shown in blue and second-length path in orange (using the reduced expressions from Example 5.8).
Compatible reflection orders exist for every reduced expression, by Proposition 2.6. Each reflection order which is Deodhar with respect to $v$ is compatible with a unique reduced expression for $v$.
Lemma 5.6.
Let $\prec$ be a reflection order which is compatible with $\mathbf{v}$. Then the map $\mathbf{u}\mapsto\operatorname{path}(\mathbf{u})$ is a bijection from distinguished subexpressions of $\mathbf{v}$ to the increasing paths in $\Gamma$ from $u^{-1}$ to $v^{-1}$. In particular, $\operatorname{path}(\mathbf{u}^{+})$ is the longest increasing path from $u^{-1}$ to $v^{-1}$.
Proof.
Let $\gamma$ be a path in $\widehat{\Gamma}$ ending at the vertex $v^{-1}$. Then $\gamma$ may be described as $x_{0}\overset{\beta_{1}}{-}x_{1}\overset{\beta_{2}}{-}\cdots\overset{\beta_{r}}{-}x_{r}=v^{-1}$, where $x_{j}=t_{\beta_{j+1}}\cdots t_{\beta_{r}}x_{r}$. For this path to have increasing edge labels, we must have $\beta_{1}\prec\cdots\prec\beta_{r}$. For each $1\leq i\leq|\Phi^{+}|$, let $\Phi_{i}(\prec)$ be the $i$th smallest positive root in the ordering by $\prec$. Let $i_{j}$ be the index such that $\beta_{j}=\Phi_{i_{j}}(\prec)$. Then $\gamma$ has increasing edge labels if and only if $i_{1}<i_{2}<\cdots<i_{r}$. Assume now that $\ell(x_{r-1})<\ell(x_{r})$, so that the last edge of $\gamma$ is in fact a directed edge of $\Gamma$. This occurs if and only if $\beta_{r}\in N(v^{-1})$, which, since $\prec$ is compatible with $\mathbf{v}$, is equivalent to $\{\beta_{1},\ldots,\beta_{r}\}\subseteq N(v^{-1})$. In this case, the element $x_{j}^{-1}=vt_{\beta_{r}}t_{\beta_{r-1}}\cdots t_{\beta_{j+1}}$ is obtained by taking the expression $\mathbf{v}$ and deleting the entries in positions $\ell(v)-i_{r}+1,\ell(v)-i_{r-1}+1,\ldots,\ell(v)-i_{j+1}+1$, by (2). In particular, when $x_{0}=u^{-1}$, then performing this process for $x_{0}^{-1}$ gives a subexpression $\mathbf{u}$ of $\mathbf{v}$. This inverts the map $\mathbf{u}\mapsto\operatorname{path}(\mathbf{u})$. Hence the paths $\gamma$ in the undirected Bruhat graph $\widehat{\Gamma}$ which begin at $u^{-1}$ and end at $v^{-1}$, with increasing edge labels, such that the last edge is in $\Gamma$, are exactly the $\operatorname{path}(\mathbf{u})$ for subexpressions $\mathbf{u}$ of $\mathbf{v}$.
We may now verify that $\operatorname{path}(\mathbf{u})$ is a path in $\Gamma$ if and only if $\mathbf{u}$ is distinguished. Indeed, assume $\operatorname{path}(\mathbf{u})$ is not a path in $\Gamma$ and let $j$ an index so that $x_{j-1}\leftarrow x_{j}$. Then $\mathbf{u}_{(\ell(v)-i_{j})}s_{\ell(v)-i_{j}}<\mathbf{u}_{(\ell(v)-i_{j})}$ (here we use the fact that $s_{\ell(v)-i_{j}}s_{\ell(v)-i_{j}+1}\cdots s_{\ell(v)}$ is a reduced expression). Hence $\mathbf{u}$ is not distinguished, since $\mathbf{u}_{(\ell(v)-i_{j}+1)}=\mathbf{u}_{(\ell(v)-i_{j})}$. Conversely, if $j$ is an index so that $\mathbf{u}_{(\ell(v)-i_{j})}s_{\ell(v)-i_{j}}<\mathbf{u}_{(\ell(v)-i_{j})}$ and $\mathbf{u}_{(\ell(v)-i_{j}+1)}=\mathbf{u}_{(\ell(v)-i_{j})}$, then $\ell(x_{j-1})>\ell(x_{j}).$ This proves the claim. ∎
Remark 5.7.
Under the bijection of Lemma 5.6, Dyer’s formula (Proposition 2.5) for the $R$-polynomial turns into Deodhar’s formula in terms of distinguished subexpressions [Deo85].
Example 5.8.
Take $W=S_{4}$, with simple roots $\alpha_{1},\alpha_{2},\alpha_{3}$. Let $u=s_{\alpha_{2}}$ and
$\mathbf{v}=s_{\alpha_{2}}s_{\alpha_{3}}s_{\alpha_{1}}s_{\alpha_{2}}s_{\alpha_{1}}=s_{1}s_{2}s_{3}s_{4}s_{5}.$
Then $u^{-1}$ and $v^{-1}$ are the permutations with the one-line notations $1324$ and $3421$, respectively. The Bruhat graph $\Gamma_{u^{-1},v^{-1}}$ is shown in Figure 4. The positive subexpression is
$\mathbf{v}$ == $s_{1}$ $s_{2}$ $s_{3}$ $s_{4}$ $s_{5}$
$\mathbf{u}^{+}$ == 11 11 11 $s_{4}$ 11
Here, the notation indicates that the partial products are given by $\mathbf{u}^{+}_{(k+1)}=\mathbf{u}^{+}_{(k)}$ for $k\neq 3$, and $\mathbf{u}^{+}_{(4)}=\mathbf{u}^{+}_{(3)}s_{4}$. In this case $J_{\mathbf{u}^{+}}=\{1,2,3,5\}$. Among these indices, only $5$ is mutable. The almost positive subexpression $\mathbf{u}^{\langle 5\rangle}$ is
$\mathbf{u}^{\langle 5\rangle}$ == $s_{1}$ 11 $s_{3}$ 11 $s_{5}$
From these two distinguished subexpressions, we produce two paths:
$\operatorname{path}(\mathbf{u}^{+})$ $:u^{-1}=s_{4}\xrightarrow{\alpha_{1}}s_{5}s_{4}\xrightarrow{\alpha_{2}}s_{5}s_{4}s_{3}\xrightarrow{\alpha_{1}+\alpha_{2}+\alpha_{3}}s_{5}s_{4}s_{3}s_{2}\xrightarrow{\alpha_{2}+\alpha_{3}}s_{5}s_{4}s_{3}s_{2}s_{1}=v^{-1}$
$\operatorname{path}(\mathbf{u^{\langle 5\rangle}})$ $:u^{-1}=s_{5}s_{3}s_{1}\xrightarrow{\alpha_{1}+\alpha_{2}}s_{5}s_{4}s_{3}s_{1}\xrightarrow{\alpha_{1}+\alpha_{2}+\alpha_{3}}s_{5}s_{4}s_{3}s_{2}s_{1}=v^{-1}.$
We have labeled the edges by their root labeling. Now, to compare with Lemma 5.6, we should determine a reflection order compatible with $\mathbf{v}$. We have
$\Phi_{1}(\mathbf{v})$ $=\alpha_{1}$
$\Phi_{2}(\mathbf{v})$ $=\alpha_{1}+\alpha_{2}$
$\Phi_{3}(\mathbf{v})$ $=\alpha_{2}$
$\Phi_{4}(\mathbf{v})$ $=\alpha_{1}+\alpha_{2}+\alpha_{3}$
$\Phi_{5}(\mathbf{v})$ $=\alpha_{2}+\alpha_{3}.$
The unique reflection order which puts these roots in order as the smallest roots is
$\alpha_{1}\prec\alpha_{1}+\alpha_{2}\prec\alpha_{2}\prec\alpha_{1}+\alpha_{2}+\alpha_{3}\prec\alpha_{2}+\alpha_{3}\prec\alpha_{3}.$
We can now verify that $\operatorname{path}(\mathbf{u}^{+})$ is the longest increasing path with respect to $\prec$, and $\operatorname{path}(\mathbf{u}^{\langle 5\rangle})$ is the unique second-length path. ∎
We can now state the result motivating the construction in Theorem 3.5. Note that each element of $J_{\mathbf{u}}$ naturally labels an edge of $\operatorname{path}(\mathbf{u})$.
Proposition 5.9.
Let $\prec$ be a reflection order compatible with $\mathbf{v}$, with longest increasing path $\gamma_{0}=\operatorname{path}(\mathbf{u^{+}})$. Then, under the correspondence between elements of $J_{\mathbf{u}^{+}}$ and edges of $\operatorname{path}(\mathbf{u}^{+})$, the frozen indices correspond to elements of $F_{u^{-1},v^{-1}}$.
Proof.
An index $a\in J_{\mathbf{u}^{+}}$ is mutable if and only if $\mathbf{u}^{\langle a\rangle}$ is an almost positive subexpression. By Lemma 5.6, this is true if and only if there is a second-length path diverging from $\gamma_{0}$ at $v^{-1}\mathbf{v}_{(a)}(\mathbf{u}^{+}_{(a)})^{-1}$. And by Lemma 3.1 this happens if and only if the subsequent edge of $\gamma_{0}$ is not in $F_{u^{-1},v^{-1}}$. ∎
Remark 5.10.
A clash in standard conventions for increasing paths and for subexpressions has resulted in the inverses appearing in Proposition 5.9. However, inversion is a poset isomorphism $u^{-1},v^{-1}]\to[u,v]$ sending $F_{u^{-1},v^{-1}}$ to a diamond-generating set of $[u,v]$.